Spitting cobras can defend themselves by squeezing muscles around their venom glands to squirt venom at an attacker. Suppose a spitting cobra rears up to a height of 0.480 m above the ground and launches venom at 3.70 m/s, directed 49.0° above the horizon. Neglecting air resistance, find the horizontal distance (in m) traveled by the venom before it hits the ground.
Answer #1

2 o 2 usoso 2 -0.48 using H = utt tat -0.48 = 3.7 Sin 69 t -149.8x+ 2.87- 4.97? t = -b + 86²490 2.8 + 2.8?+ 4* 4.940.48 4942-if any doubt plzzzzz comment first.plzzzzzzzzz like????